ISOTOPIC SELECTIVITY IN THE ELECTRONIC PREDISSOCIATION DYNAMICS OF CN − Ar_{n} CLUS-TERS

\underline{YALING CHEN}, WILLIAM G. LAWRENCE, and MICHAEL C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

CN − Ar_{n} clusters with an average size of n=100, were prepared in a supersonic expansion. Electronic excitation of the clusters to CN(B^{2}Σ^{+})-Ar_{n} was followed by electronic predissociation. In a previous low-resolution study^{a}, it was proposed that predissociation involved the ejection of free CN(A^{2}Π) from the surface of the cluster. In the present work, OODR techniques were used to demonstrate that free CN(A^{2}Π) was produced. Interestingly, although a natural isotopic mixture of CN was used (1% ^{13}CN), similar quantities of ^{12}CN and ^{13}CN fragments were detected. The product state distributions resulting from predissociation, and the isotopic selectivity of the dynamics will be discussed.

aH.-S. Lin, M. G. Erickson, Y. Lin, W. H. Basinger, W. G. Lawrence, M. C. Heaven, Chem. Phys. 189, 235 (1994)